OpenStudy (anonymous):

which angle corresponds to a circular arc on the unit circle with length of 2?

OpenStudy (eyust707):

i believe arc length equals R times the angle(in radians)

OpenStudy (dumbcow):

since radius of unit circle is 1 angle = 2 radians

OpenStudy (anonymous):

i need a degree though

OpenStudy (dumbcow):

multiply by 180/pi

OpenStudy (anonymous):

180/pi?

OpenStudy (anonymous):

the answer...

OpenStudy (dumbcow):

yeah its the conversion factor pi radians = 180 degrees

OpenStudy (anonymous):

didn't work though :/

OpenStudy (eyust707):

dumbcow may look kinda dumb...

OpenStudy (eyust707):

hehe but im not so sure he is

OpenStudy (anonymous):

let me try something real quick

OpenStudy (dumbcow):

haha 2*(180/pi) = 360/3.14 = 114.59

OpenStudy (anonymous):

good call

OpenStudy (anonymous):

thanks

OpenStudy (eyust707):

=P